# TRMAC: A Time-Reversal-based MAC Protocol for Wireless Networks within Computing Packages

### v1

@@ Abstract
As chiplet-based integration and many-core architectures become the norm in high-performance computing, on-chip wireless communication has emerged as a compelling alternative to traditional interconnects. However, scalable Medium Access Control (MAC) remains a fundamental challenge, particularly under dense traffic and limited spectral resources. This paper presents TRMAC, a novel cross-layer MAC protocol that exploits the spatial focusing capability of Time Reversal (TR) to enable multiple parallel transmissions over a shared frequency channel. By leveraging the quasi-deterministic nature of on-chip wireless channels, TRMAC pre-characterizes channel impulse responses to coordinate access using energy-based thresholds, eliminating the need for orthogonal resource allocation or centralized arbitration. Through detailed physical-layer simulation and system-level evaluation on diverse traffic, TRMAC demonstrates comparable or superior performance to existing multi-channel MAC protocols, achieving low latency, high throughput, and strong scalability across hundreds of cores. TRMAC provides a low-complexity, high-efficiency solution for future Wireless Networks-on-Chip (WNoCs), particularly in chiplet-based systems where spatial reuse and modularity are critical. With simulations we prove that TRMAC can be utilized for parallel transmissions with a single frequency channel with a similar throughput and latency as in using multiple frequency bands omitting the need for complex transceivers. This work establishes a new design direction for MAC protocols that are tightly integrated with the underlying channel physics to meet the demands of next-generation computing platforms.
@@
journal: Nano Communication Networksfn1fn1footnotetext: Authors gratefully acknowledge funding from the European Commission through projects with GA 101042080 (WINC), 101099697 (QUADRATURE) and 101189474 (EWiC).
@@ 1 Introduction
Recent advances in High-Performance Computing (HPC), driven by the proliferation of CPUs, GPUs, and domain-specific accelerators, have led to the integration of many cores within a single computing package, and more recently, the disaggregation of functionality across chiplets interconnected via high-density packaging [1]. To fully harness the computational capabilities of such architectures, whether monolithic or chiplet-based, parallel processing has become indispensable, enabling the concurrent execution of tasks and efficient handling of large-scale data. However, this parallelism introduces significant challenges in inter-core and inter-chiplet communication, particularly in maintaining synchronization and ensuring coherent data sharing [2, 3, 4]. Consequently, precise and efficient communication mechanisms are critical to coordinating tasks and maximizing system performance [5, 6, 3].
@@
Wireless Networks-on-Chip (WNoCs) have recently emerged as a promising complement to traditional wired NoC architectures, offering a path to alleviate the growing challenges of latency and power dissipation in massively multi-core and System-on-Chip (SoC)/System-in-Package (SiP) designs [5, 7]. In these systems, communication is facilitated through an interposer-based environment, where antennas integrated on each core or chiplet establish wireless links within the package. Compared to conventional wired interconnects, wireless links offer substantial advantages, including single-hop data transfers, inherent broadcasting capabilities, and dynamic reconfigurability [7, 8]. However, to be viable in high-performance environments, WNoCs must meet stringent requirements: data rates exceeding 100 Gb/s, latencies below 10 ns, energy consumption below 1 pJ/bit, and Bit Error Rates (BERs) below $10^{-10}$, ensuring performance, efficiency and reliability [9].
@@
While the use of wireless communication within chip packages offers compelling benefits, it also presents unique challenges. The confined and reflective nature of the chip environment leads to significant electromagnetic energy dispersion, resulting in high delay spread and Inter-Symbol Interference (ISI) that degrades communication at high data rates. Furthermore, simultaneous non-orthogonal transmissions over shared bandwidths introduce severe Co-Channel Interference (CCI), particularly in parallel wireless links. To ensure robust and efficient data exchange, it is crucial to adopt Medium Access Control (MAC) or multiple access schemes that begin at the physical layer to mitigate interference and avoid collisions.
@@
Figure 1: WNoC in an interposer package with two parallel communication links. Data packets are transmitted using TR filters, which suppress CCI at neighboring nodes. Spatial channels are coordinated via TRMAC.
@@
Fortunately, the quasi-deterministic nature of the on-chip wireless channel, enabled by fixed package dimensions and known material properties, allows for accurate pre-characterization of the Channel Impulse Response (CIR). This predictability makes Time Reversal (TR) an attractive technique for improving communication robustness [10]. By exploiting the spatial and temporal invariance of the channel, TR focuses energy both in space and time at the intended receiver as shown in Figure 1, thereby significantly reducing ISI and enhancing signal fidelity [10, 11, 12]. TR filtering employs the time-reversed version of the CIR to act as a spatial matched filter that concentrates energy at the target receiver. In reverberant environments with rich multipath effects, the uniqueness of each CIR results in stronger signal focusing and reduced interference. Additionally, TR minimizes CCI by exploiting the orthogonality of spatial channels between different transmitter-receiver pairs. This allows multiple parallel transmissions to occur simultaneously in both time and frequency domains, provided that the number of concurrent transmissions, which we refer to the Number of Parallel Transmissions (NPT), is carefully bounded. Although TR suppresses CCI, it does not eliminate it entirely, necessitating well-defined NPT thresholds, which remain stable due to the time-invariant nature of the channel.
Despite enabling spatial multiplexing, TR does not inherently prevent transmission collisions. Moreover, TR-based communication is susceptible to the deafness problem, wherein transmitters and receivers lack awareness of ongoing transmissions due to the absence of explicit coordination, beyond the NPT constraint. To address these limitations, this work introduces TRMAC: a cross-layer, adaptive MAC protocol specifically designed to manage TR-precoded transmissions in chip-scale wireless environments.
@@
Existing WNoC MAC protocols typically rely on resource-sharing schemes such as frequency-division, time-division, or code-division multiple access (FDMA, TDMA, or CDMA) to mitigate interference in multi-link environments [13, 14]. Some approaches adopt hierarchical hybrid wired-wireless architectures, but tend to emphasize overall system design while overlooking the effects of wireless channel impairments [15, 16, 17, 18, 4]. Table 1 summarizes recent MAC protocols in terms of architecture, multiplexing, scalability, and traffic awareness capabilities. Many of these schemes are tailored to specific architectures or applications, and do not scale well with increasing core counts or dynamic traffic. Additionally, they often incur extra overhead in area and power due to complex transceiver designs or control mechanisms, and may introduce latency through intricate scheduling or arbitration.
Table 1: Summary of the design traits of existing protocols
Work Architecture Multiplexing Scalability Traffic Awareness
Shamim et al. [4] Non-static N/A High Low
Mansoor et al. [19] Non-static TDMA Medium High
Franques et al. [20] Non-static N/A High High
Mestres et al. [21] Non-static N/A High Low
Vijayakumaran et al. [13] Static [22] CDMA High High
Matolak et al. [14] Static FDMA, TDMA High Medium
DiTomaso et al. [17] Static TDMA High High
TRMAC (this work) Non-Static Spatial High High
Despite significant progress, none of the existing MAC protocols account for communication over multiple spatial channels as enabled by TR. Most designs either assume global knowledge of ongoing transmissions or rely on orthogonal resource allocation to support parallelism. In contrast, TRMAC leverages the spatial diversity offered by TR to support simultaneous transmissions in both time and frequency. It adopts a cross-layer approach, spanning physical and link layers, tailored to the unique characteristics of quasi-deterministic on-chip wireless channels. TRMAC improves system efficiency by scaling the aggregate bandwidth with the number of spatial channels, while reducing latency to meet system-level performance demands. It also complements orthogonal multiplexing methods by offering an additional degree of freedom in spectral sharing for massively multi-core and multi-chiplet systems.
In previous work [23], we analyzed the feasibility of TRMAC with respect to parallel concurrent transmissions, where the protocol can be employed with a single frequency band and multiple parallel transmissions by reducing the area and power cost on utilizing multiple frequency bands. In this work we comprehensively analyzed TRMAC with respect to the physical layer characterization, wireless data transfer and heterogeneous traffic conditions. The main contributions of this work are summarized as follows:
@@
    We propose TRMAC, a cross-layer MAC protocol in detail that leverages TR filtering to enable parallel, low-interference wireless transmissions in chip-scale environments.
@@
    We analytically characterize the NPT that can be supported with TR filtering under a target BER, establishing design-time thresholds based on physical-layer channel properties.
@@
The rest of the paper is organized as follows: Section 2 provides background and motivation for wireless on-chip communication and TR-based design. Section 3 details the proposed TRMAC protocol and its physical-layer implementation. Section 5 presents the performance evaluation of TRMAC against existing approaches. Section 6 reviews relevant MAC protocols in the context of WNoC architectures. Finally, Section 7 summarizes the contributions and outlines directions for future work.
@@ 2 Background
2.1 Wireless Networks-on-Chip (WNoC)
@@ 3.1 Motivation
3.2 Protocol Working
The TRMAC protocol operates in three sequential phases: preamble transmission, ACK detection, and data transfer. These stages are illustrated in Figure 3, while the detailed working flowchart is presented in Figure 4.
@@ 3.2.1 Phase I: Preamble Transmission
When a packet is ready, the Transmitter (Tx) initiates communication by sending a preamble, filtered through the time-reversed CIR of the intended Receiver (Rx). The preamble contains the Tx node’s address. Upon receiving the preamble, the Rx decodes the address, selects the corresponding CIR, and returns a one-bit TR-based Acknowledgment (ACK). Due to spatial focusing, the ACK will reach only the intended Tx node. If the Rx erroneously decodes the preamble because of collision or interference, it may send an ACK to the wrong node, which will disregard it.
@@ 3.2.2 Phase II: Acknowledgment Detection
Following the preamble transmission, the Tx waits for the ACK from the Rx. Two collision cases are considered:
@@
    Case 1: No ACK received. This occurs when the Rx is unable to decode the preamble correctly due to a collision or interference, or when the Rx simultaneously initiates its own transmission, leading to deafness.
@@
    Case 2: ACK received with high energy. The Tx receives an ACK, but with energy significantly above the pre-characterized threshold, indicating a possible collision with another ACK. This is a rare but possible scenario if multiple transmitters simultaneously attempt communication on overlapping spatial channels.
In both cases, the Tx concludes that a collision has occurred and halts further transmission until recovery actions are performed.
@@ 3.2.3 Phase III: Data Transmission
If the Tx receives an ACK with energy matching the expected threshold, it proceeds to transmit the data packet to the Rx. A successful transmission is completed across all three phases.
3.3 Decisions and Assumptions
Transmitter: Initially, frequency channels ($f_{n}$) are uniformly distributed among nodes, with each channel covering a sector of $N_{\text{cores}}/f_{n}$. Transmissions are assumed to be globally synchronized. Each Tx node starts communication from Phase I and proceeds sequentially unless a collision is detected.
To detect ACKs and possible collisions, each Tx compares the received signal energy to a predefined threshold $E_{\text{th,correct}}$. If a Tx aims to communicate with multiple nodes simultaneously through symmetric spatial channels, this threshold is adjusted using pre-characterized values to reflect the expected composite energy.
Receiver: Upon receiving a preamble, the Rx uses threshold-based detection to identify the Tx. If multiple transmissions arrive simultaneously at the same Rx, the energy may superimpose, leading to incorrect preamble decoding and misidentification of the Tx. Thus the Rx may select the wrong CIR, causing a failed transmission and contributing to overall collision rate.
Additionally, if the intended Rx initiates its own transmission at the same time and frequency, self-interference prevents it from receiving the original preamble or sending an ACK, again contributing to the deafness problem.
A comprehensive illustration of the working of transmitter and receiver in the proposed TRMAC protocol is shown in Figure 4.
3.4 Collision Handling and Channel Awareness
When a collision is detected in Phase II, TRMAC employs a two-step recovery mechanism:
@@ 3.4 Collision Handling and Channel Awareness
    Step 1: Exponential Backoff. The Tx enters a wait period that grows exponentially with each consecutive collision, reducing the probability of repeated interference during retries.
@@
    Step 2: Tone-Based Signaling. A single-bit tone is broadcast over a dedicated, high-frequency channel (170 GHz, 1 GHz bandwidth) to signal the start of Phase III, as illustrated in Figure 5. If no tone is detected, the medium is assumed to be free. This mechanism allows nodes to coordinate transmission timing without additional interference or complex arbitration.
Figure 5: Timing diagram of tone-based coordination for collision recovery.

### v4

@@ Abstract
As chiplet-based integration and many-core architectures become the norm in computing, on-chip wireless communication has emerged as a compelling alternative to traditional interconnects. However, scalable Medium Access Control (MAC) remains a fundamental challenge, particularly under dense traffic and limited spectral resources. This paper presents TRMAC, a novel cross-layer MAC protocol that exploits the spatial focusing capability of Time Reversal (TR) to enable multiple parallel transmissions over a shared frequency channel. By leveraging the quasi-deterministic nature of on-chip wireless channels, TRMAC pre-characterizes channel impulse responses to coordinate access using energy-based thresholds, eliminating the need for orthogonal resource allocation or centralized arbitration. Through detailed physical-layer simulation and system-level evaluation on diverse traffic, TRMAC demonstrates comparable or superior performance to existing multi-channel MAC protocols, achieving low latency, high throughput, and strong scalability across hundreds of cores. Moreover, we prove that TRMAC can be utilized for parallel transmissions with a single frequency channel with a similar throughput and latency as in using multiple frequency bands, omitting the need for complex transceivers.
@@
journal: Nano Communication Networksfn1fn1footnotetext: Authors gratefully acknowledge funding from the European Commission via projects with GA 101042080 (WINC), 101099697 (QUADRATURE) and 101189474 (EWiC).
@@ 1 Introduction
Recent advances in computer architecture have led to the integration of many processing cores within a single computing package, and more recently, the disaggregation of functionality across chiplets interconnected via high-density packaging [1]. To fully harness the computational capabilities of such architectures, whether monolithic or chiplet-based, efficient parallel processing has become indispensable to enable the concurrent execution of tasks and efficient handling of large-scale data. However, this parallelism introduces significant challenges in the communication across cores, within or across chiplets, that is used to maintain synchronization and ensuring coherent data sharing within the system [2, 3, 4]. Consequently, precise and efficient communication mechanisms are critical for coordinating tasks and maximizing system performance [5, 6, 3].
@@
Wireless Networks-on-Chip (WNoCs) have recently emerged as a promising complement to traditional wired NoC architectures, offering a path to alleviate the growing challenges of latency and power dissipation in massive System-on-Chip (SoC) or System-in-Package (SiP) designs [5, 7]. In these systems, wireless communication can be facilitated through antennas that are integrated on each core or chiplet and that establish wireless links within the package. Compared to conventional wired interconnects, wireless links offer substantial advantages, including low latency via single-hop data transfers, inherent broadcasting capabilities, and dynamic reconfigurability [7, 8]. However, to fully exploit the benefits of wireless interconnections in high-performance environments, future WNoCs designs would target data rates up to 100 Gb/s, latencies below 10 ns, energy consumption around 1 pJ/bit, and Bit Error Rates (BERs) below $10^{-10}$ [9].
@@
While the use of wireless communication within chip packages offers compelling benefits, it also presents unique challenges. The confined and reverberant nature of the chip environment leads to significant delay spreads due to many reflections, resulting in high Inter-Symbol Interference (ISI) that degrades communication at high data rates. Furthermore, since signals spread across the entire environment, severe Co-Channel Interference (CCI) can be expected for transmissions within the same time-frequency channel.
@@
Figure 1: Illustration of the scenario considered in this paper. Wireless Network-on-Chip (WNoC) in a multi-chip package with multiple parallel communication links. Data packets are transmitted using time reversal, which may suppress co-channel interference at neighboring nodes. Spatial channels need to be coordinated via a link-level protocol, for which we propose TRMAC in this work.
@@
Fortunately, the quasi-deterministic nature of the on-chip wireless channel, enabled by fixed package dimensions and known material properties, allows to mitigate the ISI and CCI impairments via a technique called Time Reversal (TR). TR is a signal processing technique that employs the time-reversed version of the channel response between a transmitter-receiver pair as a spatial matched filter. As a result, TR concentrates the radiated energy at the vicinity of the intended receiver and around a given time instant, as shown in Figure 1, thereby significantly reducing ISI and enhancing signal fidelity [10, 11, 12]. Additionally, TR minimizes CCI by exploiting the orthogonality of spatial channels between different transmitter-receiver pairs. This allows, in principle, multiple parallel transmissions to occur simultaneously in both time and frequency domains.
The static nature of the wireless channels in this scenario makes TR a very attractive technique for improving communication robustness, since the channel response needs to be measured just once per each transmitter-receiver pair for the creation of the TR filter bank. Because the channels’ responses do not change over time, the filter bank can be reused throughout the lifetime of the computer to mitigate ISI and CCI. However, the use of TR filters needs to be carefully managed at runtime since, despite enabling spatial multiplexing, TR does not inherently prevent transmission collisions. Hence, to ensure robust and efficient data exchange, it is crucial to adopt Medium Access Control (MAC) that avoid or remedy collisions.
@@
Existing MAC protocols for WNoC typically rely on resource-sharing schemes such as frequency-division, time-division, or code-division multiple access to mitigate interference in multi-link environments [13, 14]. Some approaches adopt hierarchical hybrid wired-wireless architectures, but tend to emphasize overall system design while overlooking the effects of wireless channel impairments [15, 16, 17, 18, 4]. More recent protocols apply combinations of multiple access techniques and propose intelligent techniques to minimize collisions [4, 19, 20, 21, 13, 14, 17, 22, 23, 24], as summarized in Section 6 and more particularly in Table 2.
However, despite significant progress in WNoC, the existing MAC protocols do not account for the unique physical layer characteristics of the chip environment. In this context, TR represents a huge opportunity since it enables spatial multiplexing in such a reverberant environment without the need of antenna arrays; yet still, none of the existing MAC protocols supports it. In fact, the literature of MAC protocols supporting time reversal is extremely scarce in any environment. To the best of our knowledge, only the work of [25] explores a TR-based MAC for underwater acoustic networks. However, the underwater channel is highly time-variant and the communication is limited by the propagation latency and the use of acoustic transducers (as opposed to the static, short-range and high-frequency nature of WNoC).
In this paper, we propose TRMAC, a cross-layer adaptive MAC protocol that is specifically designed to leverage the unique characteristics of TR applied to on-chip wireless environments. TRMAC exploits, for the first time, the spatial multiplexing capabilities of TR to increase the throughput of wireless on-chip networks without requiring extra spectral resources. To that end, our protocol addresses the main challenge of TR: the deafness problem associated with the focusing ability of TR, which makes nodes unaware of other nodes’ transmissions and that can lead to collisions. Collision management is performed in a fully distributed manner, using tone signaling, exponential backoff, and a unique mid-transmission acknowledgment mechanism that also exploits TR. In so doing, TRMAC is able to scale the aggregate bandwidth with the number of spatial channels, which could be combined with orthogonal multiplexing methods to further improve performance in massively multi-core and multi-chiplet systems.
In our previous work [26], we presented the fundamental mechanisms behind TRMAC and an initial performance evaluation for different numbers of spatial channels, frequency channels, and number of antennas. In this paper, we significantly extend the work by providing a deeper description of the different phases of the protocol, potential collision situations, physical layer aspects, together with a more comprehensive evaluation that now includes traffic patterns of variable spatial hotspotness and temporal burstiness, and an extended discussion of the protocol overheads. In summary, the main contributions of this work are summarized as follows:
@@
    We propose TRMAC, a cross-layer MAC protocol that leverages TR filtering to enable parallel, low-interference wireless transmissions in chip-scale environments.
@@
    We discuss the strengths and limitations of the proposed approach and potential ways for future implementation of TRMAC in real systems.
The rest of the paper is organized as follows. Section 2 provides the background for wireless on-chip communication and TR-based design. Section 3 details the proposed TRMAC protocol and Section 4 discuss its physical-layer implementation. Section 5 presents the performance evaluation of TRMAC against existing approaches. Section 6 reviews relevant MAC protocols in the context of WNoC architectures. Section 7 provides a discussion of the overall results as well as of the limitations of the study. Finally, Section 8 summarizes the contributions and outlines directions for future work.
@@ 2 Background
2.1 Wireless Network-on-Chip (WNoC)
@@ 3 The TRMAC Protocol
To explain the protocol, we first provide an overview of its phases in Section 3.1, then clarify the main design decisions and assumptions in Section 3.2, and finally describe the collision management mechanisms in Section 3.3.
@@
Figure 4: Timing diagram of a single transmission with TRMAC.
3.1 Overview of the Protocol
The TRMAC protocol operates in three sequential phases: preamble transmission, ACK detection, and data transfer. If a node is ready to transmit and the channel is idle, we first transmit the address of the Transmitter (Tx) to the Receiver (Rx) as the preamble, using the appropriate TR filter. The Rx receives and decodes the address from the preamble, and then transmits the Acknowledgment (ACK) to the original sender using the TR filter pointed by the source address. If there is no collision, the Rx will use the correct TR filter and hence the Tx will receive the ACK. In this case, the Tx will continue with the rest of the data transmission together with a tone signal to avoid further collisions at this stage. If there was a collision in the first stage, the Rx will decode an incorrect address and the protocol will have to start over, as further explained in subsequent sections.
These three stages are illustrated in Figures 3 and 4, while the detailed working flowchart is presented in Figure 5. In all cases, it is considered that all nodes have a TR filter bank that is filled at the system startup time using calibration measurements. Since the system is static, the filter bank is valid for the entire system lifetime.
@@ 3.1.1 Phase I: Preamble Transmission
When a packet is ready and the channel is idle, the Tx node initiates communication by sending a preamble, filtered through the time-reversed CIR of the intended Rx node as illustrated in inset 1 of Figure 6. The preamble contains the Tx node’s address. Then, the protocol proceeds to the next phase (Section 3.1.2).
Figure 6: Block diagram illustrating the transmitting and reception process in TRMAC. Tx is on the left and Rx is on the right. (1) The Tx node first transmits the time reversed preamble through the TR filter bank. (2) The Rx node demodulates the preamble and decodes the source address. (3) The Rx node transmits the ACK going through the TR filter bank based on the address demodulated in the preamble. (4) The Tx node receives the ACK, after which the Tx can transmit the rest of the data.
@@ 3.1.2 Phase II: Acknowledgment Detection
Upon receiving the preamble, the Rx decodes the address of the Tx, selects the corresponding CIR on its filter bank, and sends back a TR-based ACK. Insets 2 and 3 of Figure 6 depict the reception of the preamble and the transmission of the ACK with TR. Due to spatial focusing, the ACK should reach only the intended Tx node, which is waiting for the ACK after the preamble transmission. Two situations may occur:
@@
    No Collision: If there is no collision, the ACK reaches the intended Tx node, which then proceeds to the next phase of the protocol (Section 3.1.3). This is illustrated in the inset 4 of Figure 6.
@@
    Collision: If two transmitters try to connect to the same Rx, a collision may occur. The Rx will try to decode the source address from a signal that contains the overlap of two transmissions, and probably lead to an incorrect address. Therefore, the ACK will most likely reach nodes not originally involved in the communication. Another type of collision can happen if the intended Rx is busy trying to transmit elsewhere and misses the preamble. In both collision cases, the Tx does not receive any ACK; the Tx concludes that a collision may have occurred and halts further transmission, looping back to the first phase of the protocol after an exponential backoff phase.
@@ 3.1.3 Phase III: Data Transmission
If the Tx receives an ACK that matches the instant when the preamble transmission was made (inset 4 of Figure 6), it proceeds to transmit the rest of the data packet to the Rx using the appropriate TR filter. Also, during this phase, the Tx activates its tone (see Figure 4 and Section 3.3) to avoid any other nodes starting a transmission until the transmission is complete. This ensures that collisions cannot occur in this phase.
3.2 Design Decisions
A comprehensive illustration of the operation of TRMAC from the perspective of the transmitter and receiver is shown in Figure 5. Next, we provide some additional details on the design decisions made.
@@ 3.2 Design Decisions
    Transmitter: If we wish to extend TRMAC to multiple frequency channels, one can initially assign $f_{n}$ frequency channels are uniformly distributed among nodes, with each channel covering a sector of $N_{\text{cores}}/f_{n}$. Transmissions are assumed to be globally synchronized. Each Tx node starts communication from Phase I and proceeds sequentially unless a collision is detected. To detect ACKs and possible collisions, each Tx compares the received signal energy to a predefined threshold $E_{\text{th,correct}}$. If a Tx aims to communicate with multiple nodes simultaneously through symmetric spatial channels, this threshold is adjusted using pre-calculated values to reflect the expected composite energy.
@@
    Receiver: Upon receiving a preamble, the Rx uses threshold-based detection to identify the Tx. If multiple transmissions arrive simultaneously at the same Rx, the energy may superimpose, leading to incorrect preamble decoding and misidentification of the Tx. Thus the Rx may select the wrong CIR, causing a failed transmission and contributing to overall collision rate. Additionally, if the intended Rx initiates its own transmission at the same time and frequency, self-interference prevents it from receiving the original preamble or sending an ACK, again contributing to the deafness problem.
3.3 Collision Management
In previous sections, we mention that a transmitter can start a transmission if a packet is ready and the channel is idle. For this to occur, a transmitter should not detect a busy tone or be on a backoff period. These are two collision management mechanisms that are further explained next.
  • 1.
    Tone Signaling: If there is no collision in the first two phases, we move to the phase of data transmission. In this phase, the transmission of the payload occurs, which is generally longer than the transmission of the preamble. However, other nodes may not be aware that this transmission is occurring, because TR makes the preamble and ACK exchange to only be seen by the intended Tx and Rx thanks to its spatiotemporal focusing ability. To avoid other transmitters to attempt a preamble transmission at this point, a continuous tone is broadcast over a separate, dedicated, low-bandwidth channel. This communicates that the Phase III of the protocol is on-going in one or more transmissions. If no tone is detected, the medium is assumed to be free.
  • 2.
    Exponential Backoff: To further clarify the protocol operation when a collision occurs, let us consider the transmission scenario illustrated in Figure 3, where in Phase I, node 1 and node 3 simultaneously transmit their time-reversed preamble data to node 2. Here, with the collision occurring in node 2, the respective address of the Tx node will be erroneously detected in the physical layer. Therefore, as indicated in Figure 3, following the previous collision, an erroneous ACK will be sent to a neighboring node in the system, while the Tx node will not receive any ACK. In this scenario, as the Tx node does not receive any ACK, it identifies the possible collision and follows a process of exponential backoff, i.e. enters a wait period that grows exponentially with each consecutive collision, reducing the probability of repeated interference during retries.